If the average of m numbers is y, and when x is added to the m numbers, the average of the m+1 numbers is b, then x is equal to which of the following.Please show me the steps.
A. my+b B. 2my-mb C. b-y+my D. b(m+1)-my
E. none of these

Answer by flame8855(424)

lets assume the sum of the m numbers = A
so A/m=y so A=my
when x is added to the m numbers the total sum of these number will be A+x and the number of these numbers will be m+1
so (my+x)/m+1 = b
my+x= bm+b
x= b(m+1)-my
the solution is D